OpenStudy (anonymous):

An object moves along a straight line in such a way that it is slowing down with an acceleration of -8 meters per second per second. The object has an initial velocity of 14 meters per second. Find the distance the object travels during the first 3 seconds.

OpenStudy (anonymous):

s=position v= velocity a=acceleration t=time \[s=s \iota+v \iota \Delta t+1/2a(\Delta t)^{2}\]

OpenStudy (anonymous):

s=0+(14m/s)(3)+1/2(-8)(3^2) s=42-36 s=6meters

OpenStudy (anonymous):

But where did you get the equation you plugged the values in to? Did you derive it or just look it up? And if so how?